### v2

# Comparing Hubbard parameters from linear-response theory and Hartree-Fock-based approach

Abstract
Density-functional theory with on-site $U$ and inter-site $V$ Hubbard corrections (DFT+UU+$V$) is a powerful and accurate method for predicting various properties of transition-metal compounds. However, its accuracy depends critically on the values of these Hubbard parameters. Although they can be determined empirically, first-principles methods provide a more consistent and reliable approach; yet, their results can vary, and a comprehensive comparison between methods is still lacking. Here, we present a systematic comparison of two widely used approaches for computing $U$ and $V$, namely linear-response theory (LRT) and the Hartree–Fock-based pseudohybrid functional formalism, applied to a representative set of oxides (MnO, NiO, CoO, FeO, BaTiO3, ZnO, and ZrO2). We find that for partially occupied transition-metal $d$ states, these two methods yield consistent $U$ values, but they differ for nearly empty or fully filled $d$ shells. For O-$2p$ states, LRT always predicts large $U$ values ($\sim$10 eV), whereas the pseudohybrid formalism produces system-dependent values depending on the level of localization and hybridization for the electronic states. Even larger differences are found for the inter-site $V$: the former predicts consistently small values (<1<1 eV), while the latter produces larger values ($\sim 3$ eV), reflecting its explicit dependence on relative charge redistribution. Our results show that while parallels between these two methods exist, they rely on distinct assumptions for determining $U$ and $V$, leading to variations in predictions of material properties.
INTRODUCTION
Density-functional theory (DFT) [Hohenberg:1964, Kohn:1965] underpins modern first-principles simulations in physics, chemistry, and materials science. Despite its remarkable success, progress hinges on the development of increasingly accurate exchange-correlation (xc) functionals. Standard approximations, such as the local spin-density approximation (LSDA) and spin-polarized generalized-gradient approximation ($\sigma$-GGA), are invaluable for many systems, yet they perform poorly for transition-metal (TM) and rare-earth (RE) compounds, where localized $d$ and $f$ states suffer from large self-interaction errors (SIEs) [Perdew:1981, MoriSanchez:2006]. To address these limitations, several beyond-standard approaches have been introduced. Among the most prominent are Hubbard-corrected DFT (DFT+$U$ [anisimov:1991, Liechtenstein:1995, Dudarev:1998] and its extension DFT+UU+$V$ [Campo:2010]), meta-GGA functionals (e.g., SCAN [Sun:2015] and its variants [Bartok:2019, Furness:2020]), and hybrid functionals (e.g., PBE0 [Adamo:1999] and HSE06 [Heyd:2003, Heyd:2006]). Combinations such as SCAN+$U$ have also been explored [Gautam:2018, Long:2020, Kaczkowski:2021, Artrith:2022]. Each strategy embodies a distinct philosophy: DFT+$U$ selectively corrects partially filled orbitals (typically $d$ and ff[Kulik:2006, Kulik:2008]; meta-GGA incorporates the kinetic-energy density and satisfies a broad set of exact constraints (17 in the case of SCAN); hybrid functionals mix a fraction of Fock exchange with (semi-)local functionals. While these methods improve predictive power, they each come with intrinsic trade-offs and limitations [Timrov:2022b]. Among them, Hubbard-corrected DFT has become especially popular for its ability to enhance accuracy at modest computational cost [Himmetoglu:2014], provided reliable Hubbard parameters are available. These can be obtained either semi-empirically (by fitting to experimental data or higher-level calculations such as hybrids or $GW$ [Hedin:1965]) or through systematic calibration. Examples include tuning $U$ to reproduce binary metal formation energies [Wang:2006, Kubaschewski:1993] or to match properties such as band gaps, magnetic moments, lattice parameters, and enthalpies [LeBacq:2004, Aykol:2014, Isaacs:2017].
While semi-empirical fitting of Hubbard parameters can be effective, it becomes impractical in materials discovery where experimental data are unavailable or when multiple target properties must be simultaneously reproduced, sometimes requiring advanced optimization schemes such as Bayesian approaches [Yu:2020, Tavadze:2021]. A natural alternative is to compute Hubbard parameters from first principles using methods such as constrained DFT (cDFT) [Dederichs:1984, Mcmahan:1988, Gunnarsson:1989, Hybertsen:1989, Gunnarsson:1990, Pickett:1998, Solovyev:2005, Nakamura:2006, Shishkin:2016], Hartree-Fock-based schemes [Mosey:2007, Mosey:2008, Andriotis:2010, Agapito:2015, Lee:2020, TancogneDejean:2020], or the constrained random-phase approximation (cRPA) [Springer:1998, Kotani:2000, Aryasetiawan:2004, Aryasetiawan:2006]. However, predicted values can differ substantially among these approaches [Himmetoglu:2014, Aryasetiawan:2006, Liu:2023], and systematic comparisons of their theoretical foundations remain scarce, although notable progress is beginning to emerge [Carta:2025]. Moreover, the dependence of $U$ on factors such as Hubbard projector choice, pseudopotentials, oxidation state, and base xc functional further complicates reproducibility [Kulik:2008, Shishkin:2016, Carta:2025, Mahajan:2021, Hubparam:2023]. Therefore, further work similar to Ref. Carta:2025 is needed to systematically compare and benchmark different first-principles approaches for computing Hubbard parameters.
Here, we present a comparative study of two such first-principles approaches. The first is based on cDFT formulated within linear-response theory (LRT) [Cococcioni:2005], which determines on-site $U$ and inter-site $V$ by enforcing piecewise linearity of the total energy with respect to the occupation of localized orbitals (Hubbard manifold), thereby reducing SIEs [Himmetoglu:2011] and improving predictive accuracy. Standard LRT requires supercell calculations, but its reformulation within density-functional perturbation theory (DFPT) [Timrov:2018, Timrov:2021] replaces these with monochromatic reciprocal-space perturbations in a unit cell. This substantially reduces computational overhead and has enabled applications to increasingly complex systems [Mahajan:2022, Timrov:2023, Binci:2023, Gebreyesus:2023, Haddadi:2024, Bonfa:2024, Macke:2024, Warda:2025]. The second approach for computing $U$ is the Hartree-Fock-based Agapito, Curtarolo, and Buongiorno Nardelli (ACBN0) functional [Agapito:2015] and its extension to inter-site $V$ dubbed as eACBN0 [Lee:2020, TancogneDejean:2020], which derives Hubbard parameters directly within self-consistent-field (SCF) cycles from an approximate form of the screened Hartree-Fock interaction energy. Using projections of Kohn-Sham (KS) states onto localized orbitals, $U$ and $V$ are updated at each iteration of the SCF cycle with negligible additional cost relative to standard DFT. This method has been shown to capture electronic structures and dynamical properties with good accuracy [Yang:2021, Yang:2022], and it holds promise for accurate evaluations of electron-phonon interactions [Jang:2023, Yang:2025].
The DFT+U+V approach.
In DFT+UU+$V$ [Campo:2010, Himmetoglu:2014], a corrective Hubbard energy $E_{U+V}$ is added to the standard DFT energy $E_{\mathrm{DFT}}$:
$E_{\mathrm{DFT}+U+V}=E_{\mathrm{DFT}}+E_{U+V}.$ (1)
Unlike DFT+$U$, which only accounts for on-site interactions scaled by $U$, DFT+UU+$V$ also includes inter-site interactions between an atom and its neighbors, scaled by $V$. In the simplified rotationally-invariant formulation [Dudarev:1998], the extended Hubbard correction reads [Timrov:2021]:
$E_{U+V}$ $=$ $\frac{1}{2}\sum_{I}\sum_{\sigma mm^{\prime}}U^{I}\left(\delta_{mm^{\prime}}-n^{II\sigma}_{mm^{\prime}}\right)n^{II\sigma}_{m^{\prime}m}$ (2)
$-\frac{1}{2}\sum_{I}\sum_{J(J\neq I)}^{*}\sum_{\sigma mm^{\prime}}V^{IJ}n^{IJ\sigma}_{mm^{\prime}}n^{JI\sigma}_{m^{\prime}m}\,$,
where $I$ and $J$ index atomic sites, $m$ and $m^{\prime}$ denote magnetic quantum numbers, and $U^{I}$ and $V^{IJ}$ are the effective on-site and inter-site Hubbard parameters, respectively. The asterisk denotes that, for each atom $I$, the sum over $J$ includes all neighbors up to a defined cutoff distance. The generalized occupation matrices $n^{IJ\sigma}_{mm^{\prime}}$ are obtained by projecting KS states $\psi^{\sigma}_{v,\mathbf{k}}$ onto localized orbitals $\phi^{I}_{m}(\mathbf{r})$:
[…]
where $v$ and $\sigma$ label bands and spin, $\mathbf{k}$ runs over the first Brillouin zone (BZ), and $f^{\sigma}_{v,\mathbf{k}}$ are the KS occupations. In Eq. (3), $m\in I$ and $m^{\prime}\in J$, meaning that $m$ runs over the magnetic quantum numbers associated with a fixed principal quantum number $n$ and orbital angular momentum $l$ of atom $I$, and analogously $m^{\prime}$ for atom $J$. The corresponding generalized KS equation can be written as [Campo:2010],
$\left(\hat{\mathcal{H}}_{\mathrm{KS}}^{\sigma}+\hat{\mathcal{V}}_{U+V}^{\sigma}\right)|\psi_{v,{\bf k}}^{\sigma}\rangle=\varepsilon_{v,{\bf k}}^{\sigma}|\psi_{v,{\bf k}}^{\sigma}\rangle,$ (4)
where $\hat{\mathcal{H}}_{\mathrm{KS}}^{\sigma}$ is the KS Hamiltonian of the standard DFT, $\varepsilon_{v,{\bf k}}^{\sigma}$ is the $v$-th band energy at momentum $\bf k$} and spin $\sigma$, and $\hat{\mathcal{V}}_{U+V}^{\sigma}$ is the Hubbard potential that is derived from the Hubbard energy (2) and which reads [Timrov:2021]:
$\hat{\mathcal{V}}_{U+V}^{\sigma}$ $=$ $\sum_{I}\sum_{mm^{\prime}}\frac{U^{I}}{2}\left(\delta_{mm^{\prime}}-2n^{II\sigma}_{mm^{\prime}}\right)|\phi^{I}_{m}\rangle\langle\phi^{I}_{m^{\prime}}|$ (5)
$-\sum_{I}\sum_{J(J\neq I)}^{*}\sum_{mm^{\prime}}V^{IJ}n^{IJ\sigma}_{mm^{\prime}}|\phi^{I}_{m}\rangle\langle\phi^{J}_{m^{\prime}}|.$
In Eqs. (2) and (5), the on-site $U^{I}$ and inter-site $V^{IJ}$ terms act in opposition: the on-site term favors electron localization on atomic sites and suppresses hybridization, whereas the inter-site term enhances hybridization with neighboring atoms. Accurate first-principles evaluation of $U^{I}$ and $V^{IJ}$ is therefore crucial to correctly describe the balance between localization and hybridization.
It is important to note that the computed values of $U^{I}$ and $V^{IJ}$ depend on the choice of Hubbard projector functions $\phi^{I}_{m}(\mathbf{r})$ used to define the occupation matrix in Eq. (3) and the Hubbard potential in Eq. (5[Lee:2020, KirchnerHall:2021, Mahajan:2021]. In the following, we briefly describe how Hubbard parameters are defined and computed within LRT and eACBN0. For both methods, we employ the same projector functions, namely atomic orbitals orthogonalized using the Löwdin method [Lowdin:1950, Mayer:2002], to ensure a fair and consistent comparison of the computed $U^{I}$ and $V^{IJ}$ values.
Hubbard parameters from LRT.
Within the LRT framework, Hubbard parameters are defined from the piecewise linearity condition of the total energy [Eq. (1)] with respect to the occupation of the Hubbard manifold and computed as [Cococcioni:2005, Campo:2010]:
$U^{I}=\left(\chi_{0}^{-1}-\chi^{-1}\right)_{II}\,$, (6)
and
$V^{IJ}=\left(\chi_{0}^{-1}-\chi^{-1}\right)_{IJ}\,$, (7)
where $\chi_{0}$ and $\chi$ are the bare and self-consistent response matrices, describing how the occupation matrices respond to localized electronic perturbations. The self-consistent response matrix is given by
$\chi_{IJ}=\sum_{m\sigma}\frac{dn^{I\sigma}_{mm}}{d\alpha^{J}}\,$, (8)
where $n^{I\sigma}_{mm^{\prime}}\equiv n^{II\sigma}_{mm^{\prime}}$ denotes the on-site occupation matrix, and $\alpha^{J}$ is the strength of the perturbation on the $J$th site. The bare response $\chi_{0}$ is defined analogously but evaluated before the self-consistent adjustment of the Hartree and xc potentials.
DFPT allows this calculation to be performed without supercells, by reformulating LRT in terms of monochromatic reciprocal-space perturbations in the unit cell. This significantly reduces computational cost [Timrov:2018, Timrov:2021]. The response of the occupation matrix is then computed as:
$\frac{dn^{I\sigma}_{mm^{\prime}}}{d\alpha^{J}}=\frac{1}{N_{\mathbf{q}}}\sum_{\mathbf{q}}^{N_{\mathbf{q}}}e^{i\mathbf{q}\cdot(\mathbf{R}_{l}-\mathbf{R}_{l^{\prime}})}\Delta_{\mathbf{q}}^{s^{\prime}}n^{s\sigma}_{mm^{\prime}},$ (9)
where $\mathbf{q}$ is the wavevector of the monochromatic perturbation, $N_{\mathbf{q}}$ is the total number of perturbations, and $\Delta_{\mathbf{q}}^{s^{\prime}}n^{s\sigma}_{mm^{\prime}}$ is the lattice-periodic response of the occupation matrices to a $\mathbf{q}$-specific perturbation. Here, $I\equiv(l,s)$ and $J\equiv(l^{\prime},s^{\prime})$, with $s$ and $s^{\prime}$ indexing atoms within the unit cell and $l$ and $l^{\prime}$ indexing unit cells; $\mathbf{R}_{l}$ and $\mathbf{R}_{l^{\prime}}$ are the corresponding Bravais lattice vectors. The quantities $\Delta{\mathbf{q}}^{s^{\prime}}n^{s\sigma}_{mm^{\prime}}$ are obtained from the response KS wavefunctions, which are computed by solving $\mathbf{q}$-specific Sternheimer equations. Further details on the DFPT implementation can be found in Refs. Timrov:2018, Timrov:2021. The efficiency of DFPT for calculating $U^{I}$ and $V^{IJ}$ has been demonstrated in numerous studies, with several recent highlights [Gelin:2024, Uhrin:2025, Bastonero:2025, Chang:2025, Binci:2025, dosSantos:2025].
Hubbard parameters from eACBN0.
In the eACBN0 method, the Hubbard parameters are defined by equating the corrective Hubbard energy [Eq. (2)] with the Hartree-Fock interaction energy [Agapito:2015, Lee:2020]. This leads to the following expressions:
$U^{I}$ $=$ $\frac{\sum\limits_{\{m\},\sigma\sigma^{\prime}}\bar{P}^{II,\sigma}_{mm^{\prime}}\bar{P}^{II,\sigma^{\prime}}_{m^{\prime\prime}m^{\prime\prime\prime}}(mm^{\prime}|m^{\prime\prime}m^{\prime\prime\prime})}{\sum\limits_{mm^{\prime},\sigma}{n^{II,\sigma}_{mm}n^{II,-\sigma}_{m^{\prime}m^{\prime}}}+\sum\limits_{m\neq m^{\prime},\sigma}{n^{II,\sigma}_{mm}n^{II,\sigma}_{m^{\prime}m^{\prime}}}},$ (10)
- $\frac{\sum_{\{m\},\sigma}\bar{P}^{II,\sigma}_{mm^{\prime}}\bar{P}^{II,\sigma}_{m^{\prime\prime}m^{\prime\prime\prime}}(mm^{\prime\prime\prime}|m^{\prime\prime}m^{\prime})}{\sum_{m\neq m^{\prime},\sigma}{n^{II,\sigma}_{mm}n^{II,\sigma}_{m^{\prime}m^{\prime}}}}\,$,
and
$V^{IJ}$ $=$ $\frac{1}{2}\frac{\sum\limits_{\begin{subarray}{c}mm^{\prime},\sigma\sigma^{\prime}\end{subarray}}[\bar{P}^{II,\sigma}_{mm}\bar{P}^{JJ,\sigma^{\prime}}_{m^{\prime}m^{\prime}}-\delta_{\sigma\sigma^{\prime}}\bar{P}^{IJ,\sigma}_{mm^{\prime}}\bar{P}^{JI,\sigma^{\prime}}_{m^{\prime}m}]}{\sum\limits_{\begin{subarray}{c}mm^{\prime},\sigma\sigma^{\prime}\end{subarray}}({n^{II,\sigma}_{mm}n^{JJ,\sigma^{\prime}}_{m^{\prime}m^{\prime}}}-\delta_{\sigma\sigma^{\prime}}{n^{IJ,\sigma}_{mm^{\prime}}n^{JI,\sigma^{\prime}}_{m^{\prime}m}})}$ (11)
$\times(mm|m^{\prime}m^{\prime})\,$,
where $(mm^{\prime}|m^{\prime\prime}m^{\prime\prime\prime})$ represents the bare Coulomb interaction between electrons with $m,m^{\prime}$ of atom $I$ and $m^{\prime\prime},m^{\prime\prime\prime}$ of atom $J$. The renormalized density matrix $\bar{P}^{IJ,\sigma}_{mm^{\prime}}$ is defined as
$\bar{P}^{IJ,\sigma}_{mm^{\prime}}=\sum_{v,\mathbf{k}}f^{\sigma}_{v,\mathbf{k}}\,\bar{N}^{IJ,\sigma}_{v,\mathbf{k}}\langle\psi^{\sigma}_{v,\mathbf{k}}|\phi^{J}_{m^{\prime}}\rangle\langle\phi^{I}_{m}|\psi^{\sigma}_{v,\mathbf{k}}\rangle\,$, (12)
where $\bar{N}^{IJ,\sigma}_{v,\mathbf{k}}$ is the renormalized occupation number defined as
$\bar{N}^{IJ,\sigma}_{v,\mathbf{k}}$ $=$ $\sum_{\{I\}}\sum_{m}\langle\psi^{\sigma}_{v,\mathbf{k}}|\phi^{I}_{m}\rangle\langle\phi^{I}_{m}|\psi^{\sigma}_{v,\mathbf{k}}\rangle$ (13)
$+\sum_{\{J\}}\sum_{m^{\prime}}\langle\psi^{\sigma}_{v,\mathbf{k}}|\phi^{J}_{m^{\prime}}\rangle\langle\phi^{J}_{m^{\prime}}|\psi^{\sigma}_{v,\mathbf{k}}\rangle\,$,
where $m\in I$ and $m^{\prime}\in J$, and the summations $\sum_{\{I\}}$ and $\sum_{\{J\}}$ mean that we sum up contributions coming from all equivalent atoms of the same type. For the case of $U^{I}$-only calculations ($I=J$), the second term in Eq. (13) is dropped [Lee:2020], so $\bar{N}^{II,\sigma}_{v,\mathbf{k}}$ reduces to the Mulliken charge of atom $I$ in state $\psi^{\sigma}_{v,\mathbf{k}}(\mathbf{r})$. Moreover, $\bar{N}^{IJ,\sigma}_{v,\mathbf{k}}$ can be viewed as a position-dependent mixing parameter within the Hartree-Fock approximation, encoding on-site and inter-site screening effects in the localized subspace. It is important to note that for a fixed ground-state density, the $U^{I}$ values computed in ACBN0 remain unchanged upon inclusion of inter-site $V^{IJ}$ terms in eACBN0. However, in self-consistent implementations of eACBN0, a coupling between $U^{I}$ and $V^{IJ}$ emerges because the ground state is obtained under the influence of both on-site and inter-site terms in the Hubbard potential (5), with self-consistency enforced for both the electron density and the Hubbard parameters. Further details on the eACBN0 formalism can be found in Refs. Agapito:2015, Lee:2020, TancogneDejean:2020.
RESULTS
In this section, we present results for the computed on-site $U$ and inter-site $V$ Hubbard parameters across three classes of systems: $i)$ TM oxides with partially occupied $d$ states (MnO, FeO, CoO, and NiO); ii)ii) compounds with nominally empty $d$ states (ZrO2 and BaTiO3); and $iii)$ a compound with fully occupied $d$ states (ZnO). These categories allow us to assess similarities and differences between LRT and eACBN0 under distinct electronic configurations.
Figure 1: Protocol for computing Hubbard parameters within the “self-consistent” scheme using (a) LRT (DFPT) [Timrov:2021] and (b) eACBN0 [Lee:2020]. In the LRT workflow, Hubbard parameters are evaluated self-consistently after the electronic charge density has been fully converged. In contrast, the eACBN0 scheme updates the charge density and the Hubbard parameters simultaneously until convergence is reached. In both workflows, atomic positions are fixed and no structural relaxations are performed.
Hubbard parameters can be computed at different levels of self-consistency. The simplest approach is a “one-shot” calculation, where the parameters are evaluated using the converged DFT charge density while keeping the geometry fixed. A more advanced “self-consistent” approach iteratively updates both the charge density and the corresponding $U$ and $V$ values, but still at a fixed geometry. Finally, in the “structurally self-consistent” scheme, structural relaxations (atomic positions and unit-cell parameters) are included in the self-consistent cycle [Timrov:2021, Bastonero:2025]. In the main text, we report results obtained with the “self-consistent” scheme, i.e. calculations in which the electronic degrees of freedom are converged together with the Hubbard parameters while the geometry remains fixed. In this approach, all $U$ and $V$ values are determined consistently with the self-consistent charge density [Timrov:2021, Lee:2020], as illustrated in Fig. 1. This allows us to isolate purely electronic effects from structural ones and to assess the influence of extended Hubbard interactions in LRT and eACBN0 with high internal consistency. For completeness, in the Supplemental Information (SI) we also present results obtained using the “structurally self-consistent” scheme, in which the self-consistent cycle includes full structural optimization with updated Hubbard parameters and charge density (see Fig. S1 in the SI). In this case, the crystal structure, the Hubbard parameters, and the charge density are iterated to convergence, yielding mutually consistent electronic and crystal structures. All computational details are provided in the Methods section.
We note that results obtained using the “one-shot” scheme are not reported in this work. In this scheme, the occupation matrices of Eq. (3) and Kohn-Sham wavefunctions employed to compute the Hubbard parameters (via Eqs. (6) and (7) for LRT and Eqs. (10) and (11) for eACBN0) differ from those that are obtained during the final DFT+UU+$V$ ground-state SCF calculation, which is determined by the Hubbard energy [Eq. (2)] and the Hubbard potential [Eq. (5)]. This difference between the Hubbard parameters evaluation step and the subsequent materials-property calculations makes the “one-shot” scheme not consistent. Moreover, the eACBN0 method is never used in a “one-shot” form in practice. To avoid potential ambiguity for the reader, we therefore do not include “one-shot” results in this work.
Using the computed self-consistent Hubbard parameters, we perform DFT+$U$ and DFT+UU+$V$ calculations to evaluate various material properties. The band gap serves as our primary benchmark for accuracy. For magnetic oxides we also examine magnetic moments, while for BaTiO3 we additionally analyze structural properties relevant to ferroelectricity. The results are systematically compared with experiments.
Since the band gap is central to this study, it is useful to recall that DFT is not a spectral theory. Local and semilocal xc functionals (LSDA, $\sigma$-GGA) notoriously underestimate band gaps in transition-metal oxides (TMOs) due to strong SIEs, particularly on TM ions. DFT+$U$ can alleviate this issue, but only when the correction is applied to states that define the gap edges (for insulators) or to states near the Fermi level (for spurious DFT metals) [KirchnerHall:2021]. The inter-site $V$ correction can further improve band gaps in covalent systems, but again only when applied to orbitals directly involved in gap formation. Prior work supports these points: Ref. KirchnerHall:2021 showed that DFT+$U$ with LRT-derived $U$ improves agreement with experimental band gaps compared to (semi-)local functionals, while DFT+UU+$V$ with LRT-derived parameters yields even better accuracy in TMOs with strong covalency [Mahajan:2021, Mahajan:2022, Timrov:2023, Gebreyesus:2023]. For $p$-block semiconductors without TM ions, Ref. Lee:2020 demonstrated that DFT+UU+$V$ with eACBN0-derived parameters accurately reproduces experimental band gaps. These insights provide the foundation for our analysis below.
MnO, FeO, CoO, and NiO
Figure 2: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for TMOs without on-site corrections on O-2pp states ($U_{O}$), using the workflow described in Fig. 1. (a) On-site Hubbard $U_{TM}$ parameter for 3dd orbitals of TM ions and inter-site Hubbard $V_{TM,O}$ parameter between 3dd orbitals of TM ions and 2pp orbitals of oxygen, (b) band gap, and (c) magnetic moment. LRT and eACBN0 denote the method used to obtain the Hubbard parameters. The experimental band gaps and magnetic moments for MnO [Kurmaev2008prb, Elp15301991prb, Cheetham1983prb, Fender1968JCP], FeO [Bowen1975JSSC, Roth1958prb, Battle1979JPCSSP], CoO [Elp60901991prb, Khan1970prb, Jauch2001prb], and NiO [Kurmaev2008prb, Sawatzky1984prl, Cheetham1983prb, Fender1968JCP] are indicated by the dashed blue lines in panels (b) and (c), respectively.
Figure 3: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for TMOs with on-site $U_{O}$ corrections on O-2pp states. The remainder of the caption is analogous to Fig. 2.
TMOs are well-known systems where SIEs are prominently manifested due to localized 3dd orbitals. Standard DFT calculations at the LDA or GGA level typically exhibit considerable discrepancies compared to experimental measurements, necessitating the use of the DFT+$U$ or DFT+UU+$V$ method. Particularly, 3dd electrons of TM ions exhibit significant spatial localization, thus requiring Hubbard $U$ corrections, while $V$ might be necessary to describe their hybridization with ligands. Therefore, we select representative 3dd TMOs – MnO, FeO, CoO, and NiO – as first targets and apply two Hubbard parameter calculation methods, LRT and eACBN0, to understand the trends in Hubbard parameters and their influence on materials properties such as band gaps and magnetic moments.
To compare the two methods under as equivalent conditions as possible, we perform LRT and eACBN0 calculations using the “self-consistent” scheme described in Fig. 1 and report the Hubbard parameters in Figs. 2 and 3. We focus first on the $U$ values shown as bar plots (unhatched) in Fig. 2 obtained in the framework of the DFT+$U$ method. A clear trend emerges where $U_{TM}$ increases and magnetic moment decreases with the number of 3dd electrons. Furthermore, both calculation methods do not alter the $U_{TM}$ values significantly when the inter-site interactions are included (compare the hatched and not hatched patterns in Fig. 2(a)). Focusing on MnO, the dominant Mn-$4s$ character of the lowest conduction band turns out to be insensitive to the Hubbard $U$ parameter applied to Mn-$3d$ states. The primary role of $U_{TM}$ is to enhance the splitting between occupied and unoccupied Mn-3dd states. This is confirmed by the partial density of states (PDOS), which shows that with increasing $U_{TM}$, occupied $d$ states shift to lower energies, while the unoccupied $d$ states (lying above the conduction band minimum) shift to higher energies [Yang:2022]. This mechanism is found to be general for other late TMOs like FeO, CoO, and NiO, which also exhibit a prominent TM-4ss character at the conduction band minimum at the $\Gamma$ point. For these materials, we consistently observe that while the energy splitting between occupied and unoccupied $d$ states scales proportionally with $U_{TM}$, the variation in the fundamental band gap remains minor in comparison.
When the inter-site Hubbard interactions ($V_{TM,O}$) between TM 3dd and oxygen 2pp states are included, the differences between the two methods remain almost unchanged. In Fig. 2, the results with the $V_{TM,O}$ term are illustrated with hatched patterns. LRT yields relatively small $V_{TM,O}$ values (below 1 eV), resulting in negligible differences in band gaps and magnetic moments whether $V_{TM,O}$ term is included or not in the total energy calculation. In contrast, eACBN0 consistently produces larger $V_{TM,O}$ values (close to 3 eV) across all TMOs, resulting in more noticeable changes in band gaps compared to LRT, while the magnetic moments change by a small margin. This effect arises because the $V_{TM,O}$ term facilitates electron delocalization into the overlapping regions between atomic pairs. This counteracts the strong localization favored by the $U_{TM}$ term and thus reduces the local magnetic moment, as previously shown through charge density difference analysis in our earlier studies [Yang:2022].
Figure 3 extends our analysis by including a Hubbard correction on the oxygen 2pp states ($U_{O}$), computed with both LRT and eACBN0. Within LRT, the resulting $U_{O}$ values are large ($\sim 10$ eV), consistently with previous studies [KirchnerHall:2021]. This is a consequence of the high electronegativity of oxygen: its 2pp shell is nearly fully occupied, which leads to a weak electronic response and therefore a large $U$ in the LRT framework [Gelin:2024, Yu:2014]. Despite these large $U_{O}$ values, the inter-site interaction $V_{TM,O}$ remains small (below 1 eV), indicating that TM–O hybridization has only a minor effect on the electronic structure of these TMOs even when $U_{O}$ is considered. In LRT, $U_{TM}$ and $U_{O}$ appear as diagonal elements, and $V_{TM,O}$ as off-diagonal elements of the matrix obtained by inverting the interacting and non-interacting response matrices, $\chi$ and $\chi_{0}$ [Eqs. (6) and (7)]. Including oxygen among the perturbed subspaces alters the full response matrices, because screening channels involving O are now explicitly taken into account. Once both TM and O sites are perturbed, the structure of the response matrices in DFT+$U$ and DFT+UU+$V$ becomes effectively similar: both schemes capture the same screening pathways involving TM and O states. Consequently, the value of $U_{TM}$ can change when $U_{O}$ is included in the LRT calculation. Comparing Figs. 2 and 3, we observe a slight increase of $U_{TM}$ for CoO and NiO (by 3355%), a substantial increase for FeO (18%), and almost no change for MnO. Conversely, eACBN0 yields smaller $U_{O}$ values than LRT ones, in the range of 232-3 eV. Similar to the case with the $V_{TM,O}$ correction, the converged Hubbard parameters can vary depending on the Hubbard manifolds being considered. In this self-consistent cycle, the parameters from one iteration feed back into the next as a correction term, thereby influencing the KS states and the final converged values. Comparing Figs. 2 and 3 for eACBN0, we find (consistently with LRT) that $U_{TM}$ in CoO increases slightly (by 1122%), while in FeO it slightly decreases (by 1144%) in contrast to LRT. For NiO and MnO, $U_{TM}$ exhibits some variations (small increases, decreases, or no significant change), depending on whether the DFT+$U$ or DFT+UU+$V$ scheme is employed.
The effects of $V_{TM,O}$ and $U_{O}$ on the magnetic moments are not very significant. The predicted magnetic moments for the TMOs are consistent regardless of computational schemes and they agree well with the experimental range irrespective of inclusion of $V_{TM,O}$. Conversely, the influence of $V_{TM,O}$ is pronounced on the band gap. For eACBN0, the inclusion of $V_{TM,O}$ introduces a non-negligible difference for FeO, CoO and NiO. Nevertheless, the inclusion of $U_{O}$ has a greater overall impact on the band gap magnitude than $V_{TM,O}$ does. Without $U_{O}$ (Fig. 2), both methods tend to underestimate the band gaps of MnO and NiO, while results for FeO and CoO are generally in good agreement with experiment. Conversely, including $U_{O}$ (Fig. 3) increases the band gaps, causing both methods to overestimate them for FeO and CoO. Notably, with $U_{O}$, the LRT overestimate the band gaps of FeO, CoO and NiO with and without $V_{TM,O}$. On a material-specific basis, LRT provides a band gap for MnO that is close to the experimental value, whereas eACBN0 still underestimates it. For FeO, CoO and NiO, the trend is reversed: LRT overestimates the gap, while eACBN0 is closer to experiment.
Finally, following the protocol in Fig. S1 in the SI [Timrov:2021], we compute the self-consistent Hubbard parameters and their corresponding optimized structural parameters using both LRT and eACBN0. The results are presented in Fig. S3 and S4 and their discussion is in the SI. For both methods, the structural parameters (lattice constants and rhombohedral angles) are generally in good agreement with available experimental data.
ZrO2 and BaTiO3
To further compare the LRT and eACBN0 methods, we examine ZrO2 and BaTiO3, where both Zr4+ and Ti4+ ions have nominal $d^{0}$ electronic configurations. Since these materials are nonmagnetic, the accuracy of the predictions will be assessed based on the band gap as well as the structural properties in the case of BaTiO3.
Figure 4 shows the results for ZrO2 without the $U_{O}$ correction. In LRT, the $U$ parameter for Zr-$4d$ states ($U_{Zr}$) is $\sim 3$ eV, reflecting the finite response arising from the small but nonzero occupancy of Zr-$4d$ states in the PBEsol ground state. In contrast, ACBN0, being more sensitive to the small residual occupancy of the nominal $d^{0}$ states, yields a much smaller $U_{Zr}$ value of $\sim 0.3$ eV, which is one order of magnitude smaller than the LRT value. Consequently, the very small $U_{Zr}$ obtained from ACBN0 leads to a significantly smaller band-gap correction in DFT+$U$ compared to that obtained using LRT [see Fig. 4(b)].
For the inter-site interactions between Zr-4dd and O-2pp states ($V_{Zr,O}$), LRT yields values below 1 eV, comparable to those obtained for the TMOs discussed in the previous section. Including these interactions in the DFT+UU+$V$ calculation increases the band gap by 7% relative to DFT+$U$ with only $U_{Zr}$ from LRT. In contrast, eACBN0 produces substantially larger $V_{Zr,O}$ values (2.3–2.5 eV), leading to a 22% band-gap increase compared to DFT+$U$ with only $U_{Zr}$, as shown in Fig. 4(b). Since $V_{Zr,O}$ depends on the interatomic distances between Zr and its nearest-neighbor O atoms, a range of values is reported in Fig. 4(a) rather than a single number. Despite these $V_{Zr,O}$ corrections, both methods still underestimate the experimental band gap by more than 1 eV [Bersch:2008].
Including $U_{O}$ substantially increases the band gap, as shown in Fig. 5. Both methods (LRT and eACBN0) yield large $U_{O}$ values in the range of 7–9 eV, with LRT producing the higher values and eACBN0 the lower ones. Incorporating these $U_{O}$ corrections within either DFT+$U$ or DFT+UU+$V$ consistently increases the band gap relative to calculations without $U_{O}$, bringing the results closer to the experimental value. Overall, LRT-based predictions (whether or not $V_{Zr,O}$ is included) tend to overestimate the band gap, whereas eACBN0-based predictions underestimate it within DFT+$U$ and slightly overestimate it within DFT+UU+$V$.
The fully relaxed structures of ZrO2 are well described by both methods, as shown in Fig. S4, and they are only weakly affected by the inclusion of $V_{Zr,O}$ and $U_{O}$. We also note that the method-dependent variations of the “structurally self-consistent” Hubbard parameters during structural relaxation follow the same trends discussed above for the “self-consistent” case at a fixed PBEsol geometry.
Figure 4: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZrO2 without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Zr}$ parameter for 4dd orbitals of Zr ions and inter-site Hubbard $V_{Zr,O}$ parameter between 4dd orbitals of Zr ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap [Bersch:2008] is shown by the dashed blue line in panel (b).
Next, we turn to BaTiO3, which is similar to ZrO2 in the sense that Ti4+ also has a nominal $d^{0}$ configuration. In this case, we consider not only the band gap but also the crystal-structure properties, which are highly sensitive to Hubbard corrections [Gebreyesus:2023, Choi:2025]. BaTiO3 is a well-known ferroelectric material, exhibiting a rhombohedral distortion at low temperatures. However, standard DFT+$U$ with sizable $U$ corrections for Ti-$3d$ states incorrectly stabilizes a centrosymmetric cubic structure, thereby suppressing its intrinsic polarization [Tsunoda2019prm]. However, including the inter-site interactions ($V_{Ti,O}$) between Ti-$3d$ and O-$2p$ states preserves their hybridization, which is essential for maintaining the rhombohedral distortion and thus restoring the polarization [Gebreyesus:2023, Choi:2025].
Figure 5: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZrO2 with on-site $U_{O}$ corrections on O-2pp states. The remainder of the caption is analogous to Fig. 4.
In Fig. 6 we present the results obtained using the “self-consistent” scheme described in Fig. 1 without the $U_{O}$ correction. Within LRT, the Hubbard parameter for the Ti-3dd states ($U_{Ti}$) is $\sim 6$ eV, reflecting the finite response associated with the small but nonzero 3dd occupancy in the PBEsol ground state, similarly to the behavior observed in ZrO2. Since Ti-3dd states are more localized than Zr-4dd states, the resulting $U_{Ti}$ is roughly twice as large as $U_{Zr}$ in LRT. In contrast, ACBN0 yields a very small $U_{Ti}$ of $\sim 0.3$ eV, identical to $U_{Zr}$, since the Ti-3dd and Zr-4dd occupancies are both very low. Thus, ACBN0 appears largely insensitive to the degree of $d$-state localization (3dd vs. 4dd), depending primarily on the actual orbital occupancies. As shown in Fig. 6, these differences lead to markedly different DFT+$U$ outcomes: LRT-based DFT+$U$ underestimates the band gap by $\sim 13$%, whereas eACBN0-based DFT+$U$ underestimates it by $\sim 31$%. This picture changes substantially when $V_{Ti,O}$ is included. The $V_{Ti,O}$ values obtained from LRT and eACBN0 are similar in magnitude to those found for ZrO2. With these inter-site interactions included, both methods yield band gaps in excellent agreement with experiment within DFT+UU+$V$. Thus, an accurate band gap is achieved either through a combination of a large $U_{Ti}$ and small $V_{Ti,O}$ (as in LRT) or a small $U_{Ti}$ and large $V_{Ti,O}$ (as in eACBN0).
Figure 6: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for BaTiO3 without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Ti}$ parameter for 3dd orbitals of Ti ions and inter-site Hubbard $V_{Ti,O}$ parameter between 3dd orbitals of Ti ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap [Wemple:1970] is shown by the dashed blue line in panel (b).
Figure 7: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for BaTiO3 with on-site $U_{O}$ corrections on O-2pp states. The remainder of the caption is analogous to Fig. 6.
As shown in Fig. 7(a), the $U_{O}$ values obtained from LRT and eACBN0 are comparable to those for ZrO2 (see Fig. 5(a)). Including these $U_{O}$ corrections has a strong impact on the band gap of BaTiO3. In particular, when $U$ corrections are applied to both Ti-3dd and O-2pp states (and even when the inter-site term VTi,OV{Ti,O} is also included) both methods substantially overestimate the band gap (Fig. 7(b)). The only exception is the DFT+$U$ calculation with ACBN0-derived $U$ values, which still underestimates the band gap but by a smaller margin (7%).
Figure S5 in the SI shows the results of the “structurally self-consistent” calculations for BaTiO3. The Hubbard parameters $U_{Ti}$ and $V_{Ti,O}$ are only slightly modified compared with those obtained using the PBEsol geometry (compare Fig. S5(c) with Fig. 6(a)). Without $U_{O}$, the two methods exhibit different structural trends: LRT-based DFT+$U$ predicts a cubic phase ($\alpha=90^{\circ}$), whereas eACBN0-based DFT+$U$ retains the rhombohedral distortion ($\alpha<90^{\circ}$), as shown in Fig. S5(b). When $V_{Ti,O}$ is included, both methods capture the rhombohedral distortion, but their quantitative predictions differ. Specifically, DFT+UU+$V$ based on LRT parameters yields a lattice parameter, rhombohedral angle, and band gap in very good agreement with experiment, whereas DFT+UU+$V$ based on eACBN0 parameters shows significantly larger deviations for all three properties.
When $U_{O}$ is included, the structural behavior changes, as shown in Figs. S5(e)–(h). In this case, LRT fails to capture the ferroelectric nature of BaTiO3: the large $U_{Ti}$ and $U_{O}$ values overly localizes the Ti-$3d$ and O-2pp states, respectively, driving the system into a non-polar cubic phase even when the small inter-site interaction $V_{Ti,O}$ is added (Fig. S5(f)). A similar trend occurs in eACBN0-based DFT+$U$ calculations, where the combined inclusion of small $U_{Ti}$ and large $U_{O}$ suppresses the polarization and restores cubic symmetry. However, unlike LRT, adding large $V_{Ti,O}$ within the eACBN0 framework recovers the rhombohedral distortion and the associated polarization, consistent with previous findings [Choi:2025]. Nonetheless, the resulting lattice parameter and band gap remain substantially overestimated.
These results highlight that, although both approaches rely on similar Hubbard correction schemes, they lead to qualitatively different effects on structural and electronic properties of BaTiO3. Namely, LRT yields the most accurate properties when $U_{Ti}$ and $V_{Ti,O}$ are included but $U_{O}$ is excluded, whereas eACBN0 achieves its best accuracy when all correction terms are included.
ZnO
Figure 8: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZnO without on-site corrections on O-2pp states ($U_{O}$). (a) On-site Hubbard $U_{Zn}$ parameter for 3dd orbitals of Zn ions and inter-site Hubbard $V_{Zn,O}$ parameter between 3dd orbitals of Zn ions and 2pp orbitals of oxygen, and (b) band gap. The experimental band gap [Mang1995ssc, Chen:1998, Reynolds1999prb, Dong2004prb] is shown by the dashed blue lines in panel (b).
Figure 9: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for ZnO with on-site $U_{O}$ corrections on O-2pp states. The remainder of the caption is analogous to Fig. 8.
After examining TMOs with partially filled $d$ states (from Mn2+ to Ni2+) and then early TM compounds with nominal $d^{0}$ configurations (Ti4+ and Zr4+), we now turn to a material with a late $d$-block element: ZnO, where Zn2+ has a nominally filled $d^{10}$ configuration. ZnO represents the opposite extreme of the $d$-electron occupancy spectrum relative to the $d^{0}$ systems, offering a distinctly different test case for evaluating Hubbard parameters.
We first discuss the results for ZnO without the $U_{O}$ correction. As shown in Fig. 8, the LRT method yields an unphysically large self-consistent $U_{Zn}$ exceeding 100 eV. Even without performing the full self-consistent cycle (see Fig. 1), i.e. using the “one-shot” scheme, the LRT estimate already exceeds 25 eV. This well-known limitation of LRT arises from the fully occupied Zn-3dd shell, which leads to spuriously large $U$ values [Yu:2014]. The “structurally self-consistent” scheme produces a similarly large $U_{Zn}$, as shown in Fig. S6. It is important to emphasize that in ZnO the valence-band maximum is primarily O-2pp in character, while the conduction-band minimum is dominated by Zn-4ss states. As a result, applying $U_{Zn}$ to the Zn-3dd states influences the band gap only indirectly. Indeed, LRT-based DFT+$U$ increases the PBEsol band gap from $\sim 0.7$ eV to 1.94 eV, which remains well below the experimental value of $3.3\sim 3.4$ eV [Mang1995ssc, Chen:1998, Reynolds1999prb, Dong2004prb]. In contrast, the eACBN0 approach predicts also large but more moderate $U_{Zn}$ of $\sim 16$ eV, resulting in a slightly smaller band gap of 1.52 eV, as shown in Fig. 8(b).
For the inter-site interactions between Zn-3dd and O-2pp states ($V_{Zn,O}$), LRT yields small negative values. These unphysical negative values likely originate from numerical instabilities during the inversion of response matrices whose elements are close to zero and therefore highly sensitive to numerical noise. In contrast, eACBN0 produces significantly larger and more stable $V_{Zn,O}$ values of about 3 eV, consistently with the trends observed for the other materials discussed above. However, these inter-site corrections have no effect on the band gap, as shown in Fig. 8(b). This insensitivity arises because the band gap in ZnO is defined between O-2pp (valence) and Zn-4ss (conduction) states, while the Zn-3dd states lie much deeper in energy. As a result, the $V_{Zn,O}$ term mainly alters the hybridization between the O-2pp and Zn-3dd states but does not influence the band edges, leaving the band gap unchanged.
Including $U_{O}$ in addition to $U_{Zn}$, as shown in Fig. 9, leads to markedly different outcomes. In LRT, the already unphysical $U_{Zn}$ value combined with a large $U_{O}\sim 11$ eV results in a band gap that overestimates the experimental value by 44%. In contrast, eACBN0 yields still large but more moderate values, $U_{Zn}\sim 15$ eV and $U_{O}\sim 7$ eV, which together produce a band gap in excellent agreement with experiment. As before, adding the inter-site term $V_{Zn,O}$ has no influence on the band gap for the reasons discussed above. These results show that the band gap is highly sensitive to $U_{O}$, since this correction directly shifts the O-2pp states that form the valence-band maximum in ZnO.
When structural relaxations are included, both the Hubbard parameters and the resulting band gaps follow similar trends as described above (see Fig. S6). We also note that, with these Hubbard corrections, both methods reproduce the experimental structural parameters very well (see Fig. S6(a) and (d)).
DISCUSSION
Based on the numerical results presented in the previous section, we now provide a discussion and interpretation of the observed trends in the behavior of the on-site $U$ and inter-site $V$ Hubbard parameters obtained from LRT and eACBN0. Despite the fundamentally different mathematical formulations and physical concepts underlying these two methods, there are cases where they yield comparable $U$ values, as well as cases where the results diverge dramatically. By contrast, the inter-site $V$ parameters systematically exhibit large discrepancies between the two approaches. In what follows, we analyze these similarities and differences in more detail.
We start with the on-site $U_{TM}$ values. For TMOs with partially occupied $3d$ shells, LRT and eACBN0 predict similar $U_{TM}$ values except $U$ for Fe in FeO that differs by 2233 eV. However, when the $3d$ shell is nearly empty or fully occupied, the two methods behave in fundamentally different ways. In particular, for nominally empty $3d$ states, LRT still produces finite and sizable $U_{TM}$ values in the range of 3366 eV, while for fully occupied $3d$ states it yields unphysically large values. The latter is a well-known limitation of the LRT framework [Yu:2014]: for filled shells, the occupation matrices exhibit vanishing variations under a local electronic perturbation, leading to nearly singular response matrices whose inversion generates spuriously large $U_{TM}$ values. The behavior of eACBN0 is markedly different: It produces $U_{TM}$ values that are finite and in the range of physical values even for closed shells. The root cause of this behavior lies in the definition of the auxiliary objects - the renormalized occupation numbers [Eq. (13)] and renormalized density matrices [Eq. (12)]. These quantities are constructed to ensure that the sum of the Mulliken charge of each atom and the interatomic charge reproduces the correct total electronic population. As a consequence, the Hubbard parameters derived from eACBN0 naturally become very small for nearly empty shells, while for fully filled $3d$ shells they approach values of $\sim 16$ eV. This trend is not mirrored in LRT because of the different physical idea: $U_{TM}$ is computed from the occupation-matrix variations in response to local electronic perturbations. For nearly empty shells, even an infinitesimal local perturbation can induce relatively large fractional changes in the occupation matrix, which translates into sizable $U_{TM}$ values. Conversely, for fully occupied shells, the occupation matrix is extremely rigid, and the small changes captured by the linear-response equations produce nearly vanishing matrix elements. The inversion of these nearly singular matrices amplifies the response artificially, resulting in unphysically large $U_{TM}$ values. Thus, while LRT and eACBN0 can yield comparable results for partially filled shells, their behavior diverges sharply at the two extremes of the $d$-electron occupation spectrum.
Regarding the on-site Hubbard $U_{O}$ parameter for O-$2p$ states, as obtained from LRT and eACBN0, we observe notable differences. On the one hand, LRT consistently predicts large $U_{O}$ values, close to $10$ eV, essentially independent of the specific compound under investigation. This apparent universality suggests that in LRT the variation of the occupation matrix for O-$2p$ states is very similar across materials, regardless of their degree of localization or hybridization. In contrast, the eACBN0 approach provides a much broader and more system-dependent distribution of $U_{O}$ values, ranging from $2$ to $8$ eV. Specifically, $U_{O}$ is about 2233 eV for TMOs, but increases to 7788 eV in ZnO, ZrO2 and BaTiO3. These results indicate that the $U_{O}$ value does not follow the same systematic trend as the $U_{TM}$ values, which generally scale with the electron count. This behavior can be rationalized by considering that the degree of localization and hybridization of O-$2p$ states varies strongly across materials. In some cases, such as in ZrO2 and BaTiO3, the oxygen states are more localized, resulting in larger $U_{O}$ values. Additionally, it is important to account for the fact that O-$2p$ states participate in significant charge redistribution and hybridization with TM-$3d$. This naturally calls for the inclusion of inter-site $V$ terms in order to properly capture the covalency of such interactions, since purely on-site corrections are insufficient to describe the entangled nature of oxygen states.
Next, we turn to the inter-site $V$ interactions between TM-$d$ and O-$p$ states in the studied materials. From our LRT calculations, we find that for all compounds except ZnO, $V$ values are typically around $1$ eV or smaller, while in ZnO they are small and negative. Such values are characteristic of LRT: $V$ corresponds to the off-diagonal elements of the response matrix, describing how the occupation on a given atom changes when a neighboring atom is perturbed. Since these inter-site responses are much weaker than the on-site ones (which define $U$), the resulting $V$ is generally small. In ZnO, the small negative $V$ originates from numerical instabilities caused by inverting nearly singular response matrices. In contrast, eACBN0 yields a markedly different picture: The inter-site $V$ values are much larger, often comparable in magnitude to the on-site $U$. For instance, in TMOs with partially occupied $3d$ shells and in ZnO, $V$ is 2255 times smaller than the corresponding $U_{TM}$. Even more strikingly, in ZrO2 and BaTiO3, $V$ exceeds the $U_{TM}$ values of the nominally empty $d$ states by a factor of $\sim 10$. This behavior stems from the fact that in eACBN0, the formation of a chemical bond involves charge redistribution between atoms, placing $U$ and $V$ on a more equal footing than in LRT. Indeed, Eq. (13) shows that $V^{IJ}$ scales with the sum of Mulliken charges of atoms $I$ and $J$, and is therefore not intrinsically smaller than its on-site counterpart. Moreover, since eACBN0 derives $U$ and $V$ from energy contributions that depend on the spatial distribution of electronic charges, bonding tends to lower on-site $U$ while enhancing inter-site $V$. In this framework, the ratio $U/V$ reflects the balance between atomic and bond-centered charges, and is thus directly tied to bond strength. Consequently, $V$ remains sizable relative to $U$ in all systems. By contrast, in LRT the $U$ and $V$ parameters emerge from occupation matrix responses, where the dominant contribution is always local. As a result, $U$ is much larger than $V$, with ratios often an order of magnitude.
In summary, we have carried out a systematic comparison of two widely used approaches, LRT and eACBN0, for computing on-site $U$ and inter-site $V$ Hubbard parameters across three classes of materials spanning nearly empty, partially filled, and fully occupied TM-$d$ shells. These two methods rest on fundamentally different physical pictures and mathematical frameworks, yet our analysis reveals both points of agreement and distinct discrepancies. For TMOs with partially occupied $d$ states, the two methods yield broadly consistent $U$ values, whereas for nearly empty or fully filled $d$ shells the disagreement becomes substantial, reflecting intrinsic differences in how each method treats electronic interactions. For the O-$2p$ states, LRT systematically predicts large $U$ values independent of chemistry, while eACBN0 produces a much wider spread, sensitive to localization and hybridization with TM-$d$ states. The contrast is even sharper for the inter-site $V$: LRT consistently yields small values ($\sim$1 eV or less), while eACBN0 predicts much larger values, often comparable to on-site $U$. These differences can be traced to the distinct mathematical assumptions underpinning the definitions of $U$ and $V$, and to the different physical insights each method encodes - LRT focusing on response functions, eACBN0 on energetics and charge redistribution. Overall, our study highlights that while certain parallels exist for the considered set of materials, the divergences between LRT and eACBN0 are pronounced and can affect the predicted material properties for specific cases. This work represents a first step toward a systematic understanding of how these methods compare and where they diverge. Bridging their conceptual and mathematical foundations, and testing them across a broader range of materials and chemistry, remains an important direction for future developments.
METHODS
All calculations were performed using the plane-wave pseudopotential method as implemented in the Quantum ESPRESSO (QE) distribution [Giannozzi:2009, Giannozzi:2017, Giannozzi:2020]. For the xc functional, we adopted the $\sigma$-GGA with the PBEsol parametrization [Perdew:2008], and used pseudopotentials from the GBRV library (v1.5) [Garrity:2014]. The KS wavefunctions and potentials were expanded in plane waves with kinetic-energy cutoffs of 65 and 780 Ry for the TMOs, and 60 and 600 Ry for ZnO, ZrO2, and BaTiO3. The first BZ was sampled using uniform $\Gamma$-centered $\mathbf{k}$-point meshes: $\times10$$\times$10 for TMOs, $\times12$$\times$7 for ZnO, $\times7$$\times$6 for ZrO2, and $\times8$$\times$8 for BaTiO3. Spin-orbit coupling was neglected. Geometry optimizations were carried out using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [Fletcher:1987], with convergence criteria of $10^{-4}$ Ry for the total energy, $10^{-3}$ Ry/Bohr for forces, and $0.5$ kbar for pressure in TMOs. For ZnO, ZrO2, and BaTiO3, criteria of $10^{-6}$ Ry for the total energy, $10^{-5}$ Ry/Bohr for forces, and $0.5$ kbar for pressure were adopted.
The LRT Hubbard parameters were computed using DFPT [Timrov:2018, Timrov:2021] as implemented in the HP code [Timrov:2022]. Hubbard projectors were constructed from atomic orbitals orthogonalized using the Löwdin scheme [Lowdin:1950]. Uniform $\Gamma$-centered $\mathbf{k}$- and $\mathbf{q}$-point grids were employed, with sizes of $\times10$$\times$10 and $\times6$$\times$6 for TMOs, $\times12$$\times$7 and $\times6$$\times$3 for ZnO, $\times7$$\times$6 and $\times3$$\times$3 for ZrO2, and $\times8$$\times$8 and $\times4$$\times$4 for BaTiO3, respectively.
The eACBN0 Hubbard parameters were computed using our modified in-house QE package [Lee:2020, Yang:2021]. For consistency with the LRT approach, the Hubbard projectors were constructed from atomic orbitals orthogonalized using the Löwdin scheme [Lowdin:1950]. The same $\Gamma$-centered $\mathbf{k}$-point meshes used for the electronic structure calculations were employed, and convergence thresholds of $10^{-10}$ and $10^{-15}$ Ry were applied for the total eACBN0 energy and SCF cycle, respectively.
The $U$ and $V$ parameters were obtained using two self-consistent protocols, summarized in Fig. 1 and Fig. S1 of the SI. In the main text, we report the values computed by iteratively evaluating $U$ and $V$ on the PBEsol-relaxed structures using the LRT and eACBN0 methods until convergence (the “self-consistent” scheme). In the case of ZnO, however, this iterative procedure was limited to two steps due to its tendency to diverge [Yu:2014]. For completeness, the SI also presents results from the “structurally self-consistent” scheme, in which DFT+UU+$V$ structural optimizations (including Hubbard contributions to forces and stresses [Timrov:2020b]) are performed, and the Hubbard parameters are recomputed until both the structure and the $U$ and $V$ values are converged.
DATA AVAILABILITY
The data used to produce the results of this work will be available in the Materials Cloud Archive. Our modified in-house QE code for eACBN0 calculations is available at https://github.com/KIAS-CMT/DFT-U-V.
ACKNOWLEDGEMENTS
W.Y. was supported by KIAS individual Grant (No. QP090102). I.T. acknowledges support from the Swiss National Science Foundation (Grant No. 200021-227641 and No. 200021-236507). Y.-W.S. was supported by KIAS individual Grant (No. CG031509). Computer time was provided by the Swiss National Supercomputing Centre (CSCS) under project No. s1073, lp18, s1326, and the CAC of KIAS.
Supplementary Information
Self-consistent protocol with structural relaxation
Figure S1: Protocol for computing Hubbard parameters within the “structurally self-consistent” scheme using either LRT (DFPT) [Timrov:2021] or eACBN0 [Lee:2020]. Total energy, $U$, $V$, and the structural parameters are iteratively updated until each meets the chosen convergence threshold.
The computational workflow for the “structurally self-consistent” scheme, i.e. combining Hubbard-parameter calculations (via LRT or eACBN0) with structural relaxations, is illustrated in Fig. S1. The procedure begins with an initial crystal structure and an initial guess for the Hubbard parameters, which may be set to zero. A self-consistent DFT+UU(+$V$) calculation is then performed to obtain the electronic ground state. In the case of eACBN0, the Hubbard parameters $U$ and $V$ are additionally computed as described in the main text. For clarity, we emphasize that LRT and eACBN0 are executed as two independent workflows.
Next, the crystal structure is optimized using DFT+UU(+$V$), including Hubbard contributions to forces and stresses [Timrov:2020b, Yang:2021], and employing the $U$ and $V$ values determined in the preceding step. If the updated Hubbard parameters or the relaxed geometry deviate from their previous values by more than user-defined thresholds, the cycle is repeated. This iterative loop continues until both the Hubbard parameters and the crystal structure are converged. The resulting self-consistent $U$ and $V$ values are then used in production calculations.
This protocol drives the system toward an energy extremum where the electronic and structural degrees of freedom are mutually consistent. Moreover, the same workflow can be applied to DFT+$U$ alone by simply setting $V=0$. The LRT (DFPT) workflow is fully automated in AiiDA-Hubbard, which greatly streamlines the entire procedure [Bastonero:2025].
Self-consistent Hubbard parameters including structural relaxation
MnO, FeO, CoO, and NiO.
For the four transition-metal oxides (TMOs) considered here, the relaxed geometries obtained with self-consistent Hubbard parameters show very good agreement with available experimental data, as illustrated in Figs. S2 and S3. These structural results are only weakly affected by the inclusion of the on-site Hubbard correction for oxygen ($U_{O}$), as evidenced by the close similarity between Fig. S2(a) and Fig. S3(a).
The Hubbard parameters, band gaps, and magnetic moments obtained from the “structurally self-consistent” scheme (see Fig. S1) follow the same qualitative trends reported in the main text for calculations performed without structural relaxations (Figs. 2 and 3). Therefore, we conclude that full structural relaxations combined with self-consistent $U$ and $V$ values reliably reproduce the experimental geometries of MnO, CoO, and NiO, and that the associated electronic properties remain largely unchanged upon including structural relaxations.
ZrO2 and BaTiO3.
Figure S4 summarizes the results for the nominally $d^{0}$ system ZrO2 obtained using the “structurally self-consistent” scheme (see Fig. S1). The fully relaxed structural parameters show good agreement with experimental data, as illustrated in Figs. S4(a) and (d). Consistently with results from the “self-consistent” scheme (no DFT+UU+$V$ structural relaxation) presented in Figs. 4 and 5 of the main text, the LRT approach yields a finite $U_{Zr}$ value due to the response of small but non-zero $d$ occupations, whereas ACBN0 gives a vanishing value. The largest contribution to the band-gap correction arises from $U_{O}$ and $V_{Zr,O}$ (particularly within the ACBN0 scheme) which brings the calculated band gap closer to the experimental value. In contrast, the lattice constant is influenced mainly by the $U_{Zr}$ correction, with only minor sensitivity to $U_{O}$ and $V_{Zr,O}$.
Next, we consider BaTiO3, which, like ZrO2, has a nominal $d^{0}$ configuration. The results, shown in Fig. S5, follow trends consistent with those discussed in the main text. For the structural properties, when the on-site oxygen Hubbard correction ($U_{O}$) is excluded, LRT-based DFT+UU+$V$ and eACBN0-based DFT+$U$ and DFT+UU+$V$ all predict a rhombohedral distortion with an angle below $90^{\circ}$ (see Fig. S5(b)). When $U_{O}$ is included, only the eACBN0-based DFT+UU+$V$ scheme retains this rhombohedral distortion (see Fig. S5(f)).
ZnO.
Figure S6 presents the results for ZnO, a system with a nominally filled $d^{10}$ shell, obtained using the “structurally self-consistent” scheme (see Fig. S1). In line with the findings reported in the main text (without DFT+UU+$V$ structural relaxations), the LRT approach again yields an unphysically large value of $U_{Zn}$. Nevertheless, this large $U_{Zn}$ has only a minor effect on both the lattice constants and the band gap, because the Zn-$3d$ states are fully occupied and play essentially no role in the chemical bonding or the states around the band edges.
As shown in Figs. S6(d)–(f), and consistent with Fig. 9 in the main text, including $U_{O}$ increases the band gap significantly. However, while LRT leads to a band gap that substantially overestimates the experimental value, the ACBN0 scheme yields a band gap in very good agreement with experiment.
Figure S2: DFT+$U$ and DFT+UU+$V$ “structurally self-consistent” calculations for TMOs without the oxygen Hubbard correction ($U_{O}$). (a) Lattice constant $a$ and rhombohedral angle $\alpha$, (b) on-site Hubbard parameter $U_{TM}$ for the TM-$3d$ states and inter-site parameter $V_{TM,O}$ between TM-$3d$ and O-$2p$ states, (c) band gap, and (d) magnetic moment. LRT and eACBN0 indicate the method used to obtain the Hubbard parameters. Experimental lattice constants and rhombohedral angles for MnO [Cheetham1983prb], FeO [Battle1979JPCSSP], CoO [Carey1991jmr], and NiO [Cheetham1983prb] are shown as dashed blue lines in panel (a); for FeO, the values correspond to Fe0.943O. Experimental band gaps and magnetic moments for MnO [Kurmaev2008prb, Elp15301991prb, Cheetham1983prb, Fender1968JCP], FeO [Bowen1975JSSC, Roth1958prb, Battle1979JPCSSP], CoO [Elp60901991prb, Khan1970prb, Jauch2001prb], and NiO [Kurmaev2008prb, Sawatzky1984prl, Cheetham1983prb, Fender1968JCP] are indicated by dashed blue lines in panels (c) and (d), respectively.
Figure S3: DFT+$U$ and DFT+UU+$V$ “structurally self-consistent” calculations for TMOs with the oxygen Hubbard correction ($U_{O}$). The remainder of the caption is analogous to Fig. S2.
Figure S4: DFT+$U$ and DFT+UU+$V$ “structurally self-consistent” calculations for ZrO2 without $U_{O}$ [panels (a)–(c)] and with $U_{O}$ [panels (d)–(f)]. Panels (a) and (d) show the lattice constants $a$, $b$, and $c$, with experimental values [Smith1965ActaC] indicated by dashed blue lines. Panels (b) and (e) display the on-site Hubbard parameters $U_{Zr}$ and $U_{O}$ for Zr-$4d$ and O-$2p$ states, respectively, as well as the inter-site parameter $V_{Zr,O}$ between Zr-$4d$ and O-$2p$ states. Panels (c) and (f) show the corresponding band gaps, with the experimental band gap [Bersch:2008] marked by a dashed blue line.
Figure S5: DFT+$U$ and DFT+UU+$V$ “structurally self-consistent” calculations for BaTiO3 without $U_{O}$ [panels (a)–(d)] and with $U_{O}$ [panels (e)–(h)]. Panels (a) and (e) show the lattice constant $a$, while panels (b) and (f) display the rhombohedral angle $\alpha$, with experimental values [Kwei:1993] indicated by dashed blue lines. Panels (c) and (g) present the on-site Hubbard parameters $U_{Ti}$ and $U_{O}$ for Ti-$3d$ and O-$2p$ states, respectively, as well as the inter-site parameter $V_{Ti,O}$ between Ti-$3d$ and O-$2p$ states. Panels (d) and (h) show the corresponding band gaps, with the experimental band gap [Wemple:1970] marked by a dashed blue line.
Figure S6: DFT+$U$ and DFT+UU+$V$ “structurally self-consistent” calculations for ZnO without $U_{O}$ [panels (a)–(c)] and with $U_{O}$ [panels (d)–(f)]. Panels (a) and (d) show the lattice constants $a$ and $c$. Panels (b) and (e) display the on-site Hubbard parameters $U_{Zn}$ and $U_{O}$ for Zn-$3d$ and O-$2p$ states, respectively, along with the inter-site parameter $V_{Zn,O}$ between Zn-$3d$ and O-$2p$ states. Experimental lattice constants and band gaps [Mang1995ssc, Chen:1998, Reynolds1999prb, Dong2004prb] are indicated by dashed lines in panels (a) and (d), and panels (c) and (f), respectively.